# Learning to engineer protein flexibility

### v1

@@ 1 Introduction
Proteins are highly dynamic biomolecules and the presence of flexible regions is critical for their biological function (Corbella et al., 2023; Lemay-St-Denis et al., 2022). In particular, fine-tuning conformational dynamics of loops close to the active site has been shown to be an important method for modulating substrate specificities (Romero-Rivera et al., 2022), turnover rates (Crean et al., 2021), and pH dependency (Shen et al., 2021) of enzymes. In addition, the biological function of many proteins often requires that a small molecule is transported through their structures, for example via tunnels leading to the active site, whose dynamical properties are thus critical for protein function (Jurcik et al., 2018). Modulating the flexibility of protein parts even far away from the active sites has been shown to be an efficient strategy for improving proteins (Karamitros et al., 2022). However, quantification of protein flexibility remains challenging despite recent progress in both experimental and computational approaches. Experimental methods, such as X-ray crystallography, nuclear magnetic resonance, and hydrogen-deuterium exchange coupled to mass spectroscopy, require expensive equipment, specialized biophysical expertise, and time-consuming sample preparation and analysis (Peacock & Komives, 2021). Consequently, they lack the high throughput needed for systematic protein flexibility evaluation. Computational methods provide a broad spectrum of approaches for assessing protein flexibility, from rapid coarse-grained modeling to more accurate but highly computationally intensive molecular dynamics (MD) simulations (Kmiecik et al., 2016). However, the available datasets are limited, and systematic comparisons of different methods across large protein datasets are thus lacking. More importantly, it remains unclear how to effectively integrate these computational methods into the state-of-the-art generative tools, increasingly used in protein engineering and design (Kortemme, 2024).
@@
In this work, we aim to fill these gaps by proposing a new method for designing protein flexibility. First, we perform a comprehensive comparison of methods for quantification of protein flexibility and identify relevant data for learning (Section 3). Second, we address the data scarcity issue by proposing new flexibility predictors utilizing just sequential (Flexpert-Seq) or both sequential and structural (Flexpert-3D) information on the input (Section 3.3). Importantly, our goal is to create a fast predictor, which is crucial for its integration into state-of-the-art ML-guided protein engineering pipelines. Third, we introduce a new method (Flexpert-Design) for fine-tuning a protein inverse folding model to make it steerable toward desired flexibility in specified amino acids (Section 3.4). Our results show that it is possible to predict flexibility from sequence, and that adding structural information further improves the prediction. Furthermore, using our protein flexibility predictor, we demonstrate that inverse folding models can be steered toward generating protein sequences with increased flexibility, suggesting a potential strategy for engineering of protein flexibility.
@@ 2 Related work
Common experimental methods for determining protein flexibility include X-ray crystallography, Nuclear Magnetic Resonance (NMR), and Hydrogen-Deuterium Exchange coupled to Mass Spectroscopy (HDX-MS) (Peacock & Komives, 2021). Crystallography measures flexibility based on the consistency of protein conformations in a crystalline state. The regularity of atomic positions across crystal lattice cells determines the precision of atom location, quantified by the B-factor (or temperature factor), which indicates mobility (Wlodawer et al., 2008). Details on NMR and HDX-MS are given in Appendix A.
@@
Protein flexibility can be predicted using faster, more economical in silico methods (Rueda et al., 2007). A common approach involves deriving flexibility from Molecular Dynamics (MD) simulations, which apply Newton’s laws to atoms to compute their motion over time. From these simulations, Root Mean Square Fluctuations (RMSF) can be computed per residue to quantify their flexibility as observed in the simulations. While effective, MD is time-consuming, as it must explore a wide range of protein conformations. Elastic Network Models (ENMs) offer a faster alternative by modeling proteins as a system of beads and springs, using Normal Mode Analysis to predict motions (López-Blanco & Chacón, 2016; Tirion, 1996). ENMs were refined by Bahar et al. to include Gaussian and anisotropic movements (Bahar et al., 1997; Atilgan et al., 2001) and are implemented in tools like ProDy (Bakan et al., 2011; Zhang et al., 2021). More details on ENMs are given in Appendix C. Recent machine learning (ML)-based methods, such as AlphaFold (Jumper et al., 2021) and ESMFold (Lin et al., 2022), have been adapted for flexibility prediction. Combining these models with flow matching techniques has enabled sampling of broader conformational spaces at about 10 times the speed of MD, though with reduced structural precision (Jing et al., 2024).
@@
The task of predicting a sequence that folds into a given backbone structure (protein inverse folding), has become a key area where machine learning significantly contributes to protein design, along with structure prediction. ProteinMPNN (Dauparas et al., 2022), based on the graph neural network architecture introduced by Ingraham et al. (2019), is one of the most widely used tools for this task. Recent benchmarks, such as ProteinInvBench (Gao et al., 2023c), feature newer methods like KWDesign and PiFold (Gao et al., 2023b; a), which outperform ProteinMPNN in standard metrics such as sequence recovery on datasets like CATH4.3 (Pearl et al., 2003). Despite these advancements, ProteinMPNN remains the community standard, largely due to its extensive experimental validation. This highlights the limitations of metrics like sequence recovery, which do not fully capture a model’s ability to design functional or stable proteins (Wang et al., 2023). As a result, there is growing interest in developing more comprehensive evaluation metrics and tasks that better reflect real-world protein design challenges.
@@ 3.1 Methods for quantification of protein flexibility
B-factors (PDB).
The crystallographic B-factor carries information on protein flexibility, as it describes the inherent thermal vibrations of atoms in a crystal lattice. In such an organized environment however, the natural movement of the protein is diminished, and artifacts crystal packing artifacts may lead to an underestimation of the real protein motions ((Eyal et al., 2005). In this paper, we report B-factors obtained from the Protein Data Bank (PDB) (Berman et al., 2000). Note that for a few cases from PDB, we also kept the B-factors originating from NMR experiments, whose nature is different from those obtained in crystallographic experiments (for details, see Appendix A).
@@
GNM (MD str.).
Gaussian Network Models (GNM) are a type of Elastic Network Models that model the atomic motions as isotropic, i.e., they model only the magnitudes of the displacements. To quantify flexibility, we use GNM to estimate the RMSF of each residue using the input structure from the ATLAS dataset that underwent a short MD relaxation (MD str.). For details on the computation of RMSF for GNMs, see Appendix B.
ANM (MD str.).
In contrast, Anisotropic Network Models (ANM) model the atomic motions as anisotropic, i.e. they model also the direction of the displacement. As in the case of GNM, we quantify flexibility using RMSF estimated from an MD relaxed structure (MD str.). For details on the computation of RMSF for ANMs, see Appendix B.
@@ 3.2 Comparison of protein flexibility quantification methods
Table 1: Pearson correlation coefficients of flexibility predictions obtained by PDB B-factors and by computational methods (see Section 3.1 for details of the studied flexibility quantification methods). The reported coefficients are averaged over the 1390 proteins from the ATLAS dataset. As the table is symmetric, we show only values in the upper triangle. The best result in terms of correlation to MD simulations is highlighted in bold, the second best is underlined.
MD B-factors AlphaFold2 ESMFold GNM (MD str.) ANM (MD str.)
@@ 3.3 Learning to predict protein flexibility
Since inverse folding pipelines usually have access to protein backbones on the input, we also explore the strategy of using both protein sequence and protein backbone structure on the input to predict the protein sequence flexibility as close as possible to the predictions from MD trajectories. More formally, given a dataset $\mathcal{D}_{3D}=\{(\bm{s_{i}},\bm{\chi_{i}},\bm{y_{i}})\}_{i\in D}$ of protein sequences $\bm{s_{i}}\in\mathcal{A}^{N_{i}}$, protein backbone structures $\bm{\chi_{i}}\in\mathbb{R}^{4\times 3N_{i}}$ (4 atoms per residue backbone) and per residue flexibility labels $\bm{y_{i}}\in\mathbb{R}^{N_{i}}$, where $D$ is the set indexing the protein in the dataset, $\mathcal{A}$ is the alphabet of 20 natural amino acids and $N_{i}$ is the number of amino acids of protein $i$, the goal is to train a regression model $\mathcal{F}_{3D}:\mathcal{A}^{N_{i}}\times\mathbb{R}^{4\times 3N_{i}}\xrightarrow{}\mathbb{R}^{N_{i}}$ to predict the per residue flexibility from protein sequence and backbone structure. As ANMs can be run relatively fast if the protein backbone is available and they were the best performing method as evaluated in Section 3.2, we tackle the task by combining our sequence-based method Flexpert-Seq with ANM-based flexibility values. This strategy is thus akin to learning how to correct the crude flexibility annotations returned by ANMs to match more accurate values derived from MD trajectories. To this end, we enhance the architecture of Flexpert-Seq by a CNN head acting as an adaptor of the ANM-predicted flexibility into the embedding space of Flexpert-Seq (see Figure LABEL:fig:enmmethod). We train the CNN Adaptor together with the LoRA layers fine-tuning the ProtTrans Encoder of Flexpert-Seq, and together with the regression layer, mapping the combined embeddings to per-residue flexibility.
@@ 3.4 Flexpert-Design: Flexibility engineering via flexibility-aware inverse folding
More formally, we are given a dataset $\mathcal{D}_{bb}=\{(\bm{s_{i}},\bm{\chi_{i}})\}_{i\in D}$ of protein sequences $\bm{s_{i}}\in\mathcal{A}^{N_{i}}$ and protein backbone atom coordinates $\bm{\chi_{i}}\in\mathbb{R}^{4\times 3N_{i}}$ (4 atoms per residue backbone), where $D$ is the set indexing the protein in the dataset, $\mathcal{A}$ is the alphabet of 20 natural amino acids and $N_{i}$ is the number of amino acids of protein $i$. We are also given a set of flexibility instructions $F=\{\bm{f_{i}}\}_{i\in D}$, where $\bm{f_{i}}\in\mathbb{R}^{N_{i}}$. The goal is to train model $\mathcal{P}_{F}:\mathbb{R}^{N}\times\mathbb{R}^{4\times 3N}\xrightarrow{}\mathcal{A}^{N}$, which takes $N$ flexibility instructions and a backbone structure of a protein of length $N$ and maps it to a protein sequence $N$ respecting the input backbone and the flexibility instructions.
@@
where $\mathcal{L}$ is a regression loss (MSE or L1), $\bm{\chi}$ is the input backbone structure, and $\bm{\hat{s}}=\mathcal{P}_{F}(\bm{f},\bm{\chi})$ is the sequence predicted by the flexibility-aware inverse folding model $\mathcal{P}_{F}$. We illustrate our method, including the loss function $\mathcal{L}_{Flex}$, in Figure 2. Using the pseudolabels of flexibility $F^{native}$ obtained by Flexpert-3D as the flexibility instructions in training, the $L_{Flex}$ loss ensures in the training process that the model learns to preserve flexibility it has seen on the input. In inference, the learned flexibility preservation is exploited to pass non-native flexibility instructions to the model to steer the inverse folding toward generating sequences with the required flexibility.
Figure 2: Overview of the Flexpert-Design method for fine-tuning of inverse folding models toward steerability by requirements on protein flexibility. The inverse folding model ProteinMPNN, taking protein backbone structure on the input, is modified to additionally take on the input flexibility instructions. This is implemented by adding a flexibility instruction to the zero-initialized node feature for each residue. The predicted sequence is sampled in a differentiable way using Gumbel-Softmax and passed to Flexpert-3D together with the input structure. Note that we are making use of the assumption that ProteinMPNN does not modify the backbone structure as we are matching together the newly predicted sequence with the ground truth backbone structure. The flexibility of the sequence predicted by Flexpert-3D is passed to the loss function $L_{Flex}$ which computes the distance between the input flexibility instructions and the flexibility of the predicted sequence.
@@ Results.
Our predictors Flexpert-Seq and Flexpert-3D outperform the baselines when using only the sequence on the input and when also using the structure, respectively. See Table 2 for the sequence case and Table 3 for the case when using sequence and structure. Note that our predictor Flexpert-3D, which partly relies on ANM, improves over them significantly and does not rely on them exclusively (see PCC of 0.81 to ANM).
Table 2: Pearson correlation coefficients between fast sequence-based baselines (AlphaFold2 and ESMFold) and our method Flexpert-Seq (columns) and the RMSF values from MD simulations and from Anisotropic Network Models (ANM) evaluated over MD relaxed structures (rows). Note that an upper bound on correlation to MD is estimated to be 0.88. Our approach outperforms the AlphaFold2 and ESMFold baselines and obtains the best correlation with the flexibility estimates obtained from Molecular Dynamics (see the value in bold). Interestingly, the correlation coefficient of our approach and Anisotropic Network Models (ANM) is just 0.66, suggesting that both our approach and ANM might be complementary, motivating the development of a predictor combining the sequence-based predictor with a ANM.
@@
MD 0.71 0.57 0.76
ANM (MD str.) 0.56 0.43 0.66
Table 3: Pearson correlation coefficients between fast structure-based baselines (Anisotropic Network Models and Gaussian Network Models) and our method Flexpert-3D (columns) and the RMSF values from MD simulations and from Anisotropic Network Models (ANM) evaluated over MD relaxed structures (rows). Note that an upper bound on correlation to MD is estimated to be 0.88. Also note please, that our approach outperforms the ANM and GNM baselines and obtains the best correlation with the flexibility estimates obtained from Molecular Dynamics (row MD). Note that the correlation coefficient of our model with ANM is 0.81 (see the bottom row), indicating that our model relies on ANM input only partially and actually learns novel information from the sequence.
ANM (MD str.) GNM (MD str.) Flexpert-3D (ours)
MD 0.77 0.76 0.82
ANM (MD str.) 1.00 0.94 0.81

### v2

@@ 1 Introduction
Proteins are highly dynamic biomolecules, and the presence of flexible regions is critical for their biological function (Corbella et al., 2023; Lemay-St-Denis et al., 2022). In particular, fine-tuning conformational dynamics of loops close to the active site has been shown to be an important method for modulating substrate specificities (Romero-Rivera et al., 2022), turnover rates (Crean et al., 2021), and pH dependency (Shen et al., 2021) of enzymes. In addition, the biological function of many proteins often requires that a small molecule is transported through their structures, for example via tunnels leading to the active site, whose dynamical properties are thus critical for protein function (Jurcik et al., 2018). Modulating the flexibility of protein parts even far away from the active sites has been shown to be an efficient strategy for improving proteins (Karamitros et al., 2022). However, quantification of protein flexibility remains challenging despite recent progress in both experimental and computational approaches. Experimental methods, such as X-ray crystallography, nuclear magnetic resonance, and hydrogen-deuterium exchange coupled to mass spectroscopy, require expensive equipment, specialized biophysical expertise, and time-consuming sample preparation and analysis (Peacock & Komives, 2021). Consequently, they lack the high throughput needed for systematic protein flexibility evaluation. Computational methods provide a broad spectrum of approaches for assessing protein flexibility, from rapid coarse-grained modeling to more accurate but highly computationally intensive molecular dynamics (MD) simulations (Kmiecik et al., 2016). However, the available datasets are limited, and systematic comparisons of different methods across large protein datasets are thus lacking. More importantly, it remains unclear how to effectively integrate these computational methods into the state-of-the-art generative tools, increasingly used in protein engineering and design (Kortemme, 2024).
@@
In this work, we aim to fill these gaps by proposing a new method for designing protein flexibility. First, we perform a comprehensive comparison of methods for quantification of protein flexibility and identify relevant data for learning (Section 3). Second, we address the data scarcity issue by proposing new flexibility predictors utilizing just sequential (Flexpert-Seq) or both sequential and structural (Flexpert-3D) information on the input (Section 4.1). Importantly, our goal is to create a fast predictor, which is crucial for its integration into state-of-the-art ML-guided protein engineering pipelines. Third, we introduce a new method (Flexpert-Design) for fine-tuning a protein inverse folding model to make it steerable toward desired flexibility in specified positions (Section 4.2). Our results show that it is possible to predict flexibility from sequence, and that adding structural information further improves the prediction. Furthermore, using our protein flexibility predictor, we demonstrate that inverse folding models can be steered toward generating protein sequences with increased flexibility, suggesting a potential strategy for engineering of protein flexibility.
@@ 2 Related work
Common experimental methods for determining protein flexibility include X-ray crystallography, Nuclear Magnetic Resonance (NMR), and Hydrogen-Deuterium Exchange coupled to Mass Spectroscopy (HDX-MS) (Peacock & Komives, 2021). X-ray crystallography provides flexibility estimates based on the consistency of protein conformations in a crystalline state. The regularity of atomic positions across crystal lattice cells determines the precision of atom location, quantified by the B-factor (or temperature factor), which indicates mobility (Wlodawer et al., 2008). Details on NMR and HDX-MS are given in Appendix A.
@@
Protein flexibility can be predicted using faster, more economical in silico methods (Rueda et al., 2007). A common approach involves deriving flexibility from Molecular Dynamics (MD) simulations, which apply Newton’s laws to atoms to compute their motion over time. From these simulations, Root Mean Square Fluctuations (RMSF) can be computed per residue to quantify their flexibility as observed in the simulations. While effective, MD is time-consuming, as it must explore a wide range of protein conformations. Elastic Network Models (ENMs) offer a faster alternative by modeling proteins as a system of beads and springs, using Normal Mode Analysis to predict motions (López-Blanco & Chacón, 2016; Tirion, 1996). ENMs were refined by Bahar et al. to include Gaussian and anisotropic movements (Bahar et al., 1997; Atilgan et al., 2001) and are implemented in tools like ProDy (Bakan et al., 2011; Zhang et al., 2021). More details on ENMs are given in Appendix C. Recent machine learning (ML)-based methods, such as AlphaFold (Jumper et al., 2021) and ESMFold (Lin et al., 2022), have been adapted for flexibility prediction. Combining these models with flow matching techniques in AlphaFlow has enabled sampling of broader conformational spaces at about 10 times the speed of MD, though with reduced structural precision (Jing et al., 2024).
@@
The task of predicting a sequence that folds into a given backbone structure (protein inverse folding) has become a key area where machine learning significantly contributes to protein design, along with structure prediction. ProteinMPNN (Dauparas et al., 2022), based on the graph neural network architecture introduced by Ingraham et al. (2019), is one of the most widely used tools for this task. Recent benchmarks, such as ProteinInvBench (Gao et al., 2023c), feature newer methods, e.g., KWDesign and PiFold (Gao et al., 2023b; a), which outperform ProteinMPNN in standard metrics such as sequence recovery on datasets like CATH4.3 (Pearl et al., 2003). Despite these advancements, ProteinMPNN remains the community standard, largely due to its extensive experimental validation. This highlights the limitations of such metrics as sequence recovery, which do not fully capture a model’s ability to design functional or stable proteins (Wang et al., 2023). As a result, there is growing interest in developing more comprehensive evaluation metrics and tasks that better reflect real-world protein design challenges.
@@ 3.1 Methods for quantification of protein flexibility
B-factors.
The crystallographic B-factor carries information on protein flexibility, as it describes the inherent thermal vibrations of atoms in a crystal lattice. In such an organized environment, however, the natural movement of the protein is diminished, and crystal packing artifacts may lead to an underestimation of the real protein motions (Eyal et al., 2005). In this paper, we report B-factors obtained from the Protein Data Bank (PDB) (Berman et al., 2000). Note that for a few cases from PDB, we also kept the B-factors originating from NMR experiments, whose nature is different from those obtained in crystallographic experiments (for details, see Appendix A).
@@
GNM.
Gaussian Network Models (GNM) are a type of Elastic Network Models that model the atomic motions as isotropic, i.e., they model only the magnitudes of the displacements. To quantify flexibility, we use GNM to estimate the RMSF of each residue using the input structure from the ATLAS dataset that underwent a short MD relaxation. For details on the computation of RMSF for GNMs, see Appendix B.
ANM.
In contrast, Anisotropic Network Models (ANM) model the atomic motions as anisotropic, i.e. they model also the direction of the displacement. As in the case of GNM, we quantify flexibility using RMSF estimated from an MD relaxed structure. For details on the computation of RMSF for ANMs, see Appendix B.
@@ 3.2 Comparison of protein flexibility quantification methods
Table 1: Pearson correlation coefficients of flexibility predictions obtained by PDB B-factors and by computational methods (see Section 3.1 for details of the studied flexibility quantification methods). The reported coefficients are averaged over the 1383/1390 proteins from the ATLAS dataset (some were skipped due to missing pieces of structure resulting in NaNs from ENMs). As the table is symmetric, we show only values in the upper triangle. The best result in terms of correlation to MD simulations is highlighted in bold, and the second best is underlined.
MD B-factors AlphaFold2 ESMFold GNM ANM
@@ 4.1 Learning to predict protein flexibility
Since inverse folding pipelines usually have access to protein backbones on the input, we also explore the strategy of using both protein sequence and protein backbone structure on the input to predict the protein sequence flexibility as close as possible to the predictions from MD trajectories. More formally, given a dataset $\mathcal{D}_{3D}=\{(\bm{s_{i}},\bm{\chi_{i}},\bm{y_{i}})\}_{i\in D}$ of protein sequences $\bm{s_{i}}\in\mathcal{A}^{N_{i}}$, protein backbone structures $\bm{\chi_{i}}\in\mathbb{R}^{4\times 3N_{i}}$ (4 atoms per residue backbone) and per residue flexibility labels $\bm{y_{i}}\in\mathbb{R}^{N_{i}}$, where $D$ is the set indexing the protein in the dataset, $\mathcal{A}$ is the alphabet of 20 natural amino acids and $N_{i}$ is the number of amino acids of protein $i$, the goal is to train a regression model $\mathcal{F}_{3D}:\mathcal{A}^{N_{i}}\times\mathbb{R}^{4\times 3N_{i}}\xrightarrow{}\mathbb{R}^{N_{i}}$ to predict the per residue flexibility from protein sequence and backbone structure. As ANMs can be run relatively fast if the protein backbone is available and they were the best performing method as evaluated in Section 3.2, we tackle the task by combining our sequence-based method Flexpert-Seq with ANM-based flexibility values. This strategy is thus akin to learning how to correct the crude flexibility annotations returned by ANMs to match more accurate values derived from MD trajectories. To this end, we enhance the architecture of Flexpert-Seq by a CNN head acting as an adaptor of the ANM-predicted flexibility into the embedding space of Flexpert-Seq (see Figure LABEL:fig:enmmethod). We train the CNN Adaptor together with the LoRA layers by fine-tuning the ProtTrans Encoder of Flexpert-Seq and, together with the regression layer, mapping the combined embeddings to per-residue flexibility.
@@ 4.2 Flexpert-Design: Flexibility engineering via flexibility-aware inverse folding
More formally, we are given a dataset $\mathcal{D}_{bb}=\{(\bm{s_{i}},\bm{\chi_{i}})\}_{i\in D}$ of protein sequences $\bm{s_{i}}\in\mathcal{A}^{N_{i}}$ and protein backbone atom coordinates $\bm{\chi_{i}}\in\mathbb{R}^{4\times 3N_{i}}$ (4 atoms per residue backbone), where $D$ is the set indexing the protein in the dataset, $\mathcal{A}$ is the alphabet of 20 natural amino acids and $N_{i}$ is the number of amino acids of protein $i$. We are also given a set of flexibility instructions $F=\{\bm{f_{i}}\}_{i\in D}$, where $\bm{f_{i}}\in\mathbb{R}^{N_{i}}$. The goal is to train model $\mathcal{P}_{F}:\mathbb{R}^{N}\times\mathbb{R}^{4\times 3N}\xrightarrow{}\mathcal{A}^{N}$, which takes $N$ flexibility instructions and a backbone structure of a protein of length $N$ and maps it to a protein sequence of the length $N$, respecting the input backbone and the flexibility instructions.
@@
where $\mathcal{L}$ is a regression loss (MSE or L1), $\bm{\chi}$ is the input backbone structure, and $\bm{\hat{s}}=\mathcal{P}_{F}(\bm{f},\bm{\chi})$ is the sequence predicted by the flexibility-aware inverse folding model $\mathcal{P}_{F}$. We illustrate our method, including the loss function $\mathcal{L}_{Flex}$, in Figure 2. Using the pseudolabels of flexibility $F^{native}$ obtained by Flexpert-3D as the flexibility instructions in training, the $L_{Flex}$ loss ensures in the training process that the model learns to preserve flexibility it has seen on the input. During inference, the learned flexibility preservation is exploited to pass non-native flexibility instructions to the model to steer the inverse folding toward generating sequences with the required flexibility.
Figure 2: Overview of the Flexpert-Design method for fine-tuning of inverse folding models toward steerability by requirements on protein flexibility. The inverse folding model ProteinMPNN, taking protein backbone structure on the input, is modified to additionally take on the input flexibility instructions. This is implemented by adding a flexibility instruction to the zero-initialized node feature for each residue. The predicted sequence is sampled in a differentiable way using Gumbel-Softmax and passed to Flexpert-3D together with the input structure. Note that we are making use of the assumption that ProteinMPNN does not modify the backbone structure as we are matching together the newly predicted sequence with the ground truth backbone structure. The flexibility of the sequence predicted by Flexpert-3D is passed to the loss function $L_{Flex}$ which computes the distance between the input flexibility instructions and the flexibility of the predicted sequence. Icons of snowflake and flame denote frozen and unfrozen layers, respectively, in modules containing trainable parameters. Trainable (orange) and fixed (blue) modules denote whether any trainable unfrozen parameters are present in the module.
@@ Results.
Our predictors Flexpert-Seq and Flexpert-3D outperform the baselines when using only the sequence on the input and when also using the structure, respectively. See Table 2 for the sequence case and Table 3 for the case when using sequence and structure. Note that our predictor Flexpert-3D, which partly relies on ANM, improves over them significantly and does not rely on them exclusively (see PCC of 0.78 to ANM).
Table 2: Pearson correlation coefficients between fast sequence-based baselines (AlphaFold2 and ESMFold) and our method Flexpert-Seq (columns) and the RMSF values from MD simulations and from Anisotropic Network Models (ANM) evaluated over MD relaxed structures (rows). Note that an upper bound on correlation to MD is estimated to be 0.88. Evaluation performed for 139 proteins from ATLAS testset. Our approach outperforms the AlphaFold2 and ESMFold baselines and obtains the best correlation with the flexibility estimates obtained from Molecular Dynamics (see the value in bold). Interestingly, the correlation coefficient of our approach and Anisotropic Network Models (ANM) is just 0.66, suggesting that both our approach and ANM might be complementary, motivating the development of a predictor combining the sequence-based predictor with ANM.
@@
MD 0.72 0.58 0.78
ANM 0.56 0.43 0.66
Table 3: Pearson correlation coefficients between fast structure-based baselines (Anisotropic Network Models and Gaussian Network Models) and our method Flexpert-3D (columns) and the RMSF values from MD simulations and from Anisotropic Network Models (ANM) evaluated over MD relaxed structures (rows). Evaluation is performed for 139 proteins from ATLAS test set. Note that an upper bound on correlation to MD is estimated to be 0.88. Also, note that our approach outperforms the ANM and GNM baselines and obtains the best correlation with the flexibility estimates obtained from Molecular Dynamics (row MD). The correlation coefficient of our model with ANM is 0.78 (see the bottom row), indicating that our model relies on ANM input only partially and actually learns novel information from the sequence.
ANM GNM Flexpert-3D (ours)
MD 0.76 0.76 0.83
ANM 1.00 0.93 0.78